# Impact of Large Language Model Assistance on Patients Reading Clinical Notes: A Mixed-Methods Study

### v1

@@ Results
For our interview study, all seven participants had accessed a portal for patients previously, responded to a question probing their general health literacy in such a way as to indicate adequate literacy, and all but one participant agreed with the statement that AI can have a beneficial impact in medicine.
Quantitative (N=200)
Qualitative (N=7)
Highest Level of schooling
8th grade or less
3
0
High school graduate or GED
55
2
2-year college degree graduate
35
0
4-year college degree graduate
80
3
Masters or doctoral degree graduate
27
2
Race
White
158
6
Black/African American
24
0
Asian
5
1
Native American/Alaska Native
1
0
Other
4
0
More than 1
7
0
Hispanic, Latino, or Spanish origin
Yes
13
0
No
187
7
Gender
Female
196
7
Male
3
0
Other
1
0
Age
40-45
42
0
46-55
68
3
56-65
61
3
66-75
28
1
$>$76
1
0
Employment Status
Employed for wages
81
2
Self-employed
50
1
Homemaker
10
0
Unemployed
14
1
Retired
30
1
Unable to work
15
2
Accessed Portal for Notes?
Yes
122
7
No
78
0
# Notes Last Year
None
24
0
One
25
1
2 or 3
41
2
4 or more
26
4
Don’t know
6
0
General Health Literacy [26]
Adequate
167
7
Inadequate
33
0
Agreement with “Artificial Intelligence can have a beneficial impact in medicine”
Strongly Agree
57
2
Agree
100
4
Neither agree nor disagree
32
1
Disagree
4
0
Strongly Disagree
7
0
Table 1: Mixed methods study population breakdown by demographics, experience with patient portals and notes, health literacy, and aritficial intelligence sentiment.
Error Analysis
Figure 3: Error analysis of ChatGPT augmentations. (a) (b) and (c) show the number of errors in different categories, as determined by clinical reviewers. In (a), we have types of errors, such as hallucinations. The most frequent error was when ChatGPT assumed a certain context, e.g., assuming that a definition is specific to breast cancer when it truly is not. Red bars represent errors that our clinical reviewers characterized as more serious, as they could mislead a patient. In (b), this is grouped by augmented clinical notes, where P represents a participant’s donated note, and S represents a synthetic note. There are very few errors for synthetic notes, and a great deal of variation in the number of errors in donated clinical notes. In (c), the number of errors are shown in each of the augmentations, with definitions containing the most errors. (d) and (e) are the result of automated analysis of ChatGPT augmentations. (d) contains the percentage of words that are found in the Carnegie Mellon Pronouncing Dictionary Corpus Reader. Control has the least, as there is potentially more acronyms and jargon. (e) shows the Flesch Kincaid grade level score, which is generally comparable or less than control in the augmentations, except for definitions. From interviews, we know that definitions still does contain jargon that can be confusing.
We report the results of the error analysis in Fig. 3. We find that ChatGPT can make misleading errors in the context of breast cancer clinical note augmentations and identified 82 total errors. In total, 28 (34%) errors were considered to be potentially harmful and misleading (Fig. 3 (a)). However, minor errors were most common, especially assumed context (Fig. 3 (a)). For example, a definition augmentation specifying that lower outer quadrant referred specifically to the breast, though the term can refer to a specific part of other tissues as well. Many of these errors appeared in the definitions augmentation, and errors were especially common in one note’s augmentations (Fig. 3 (b-c)). In particular, most errors occurred in augmentations to real notes, while only one error appeared in augmentations to the synthetic notes (Fig. 3 (b)).
In the automated analysis, we found that the percentage of words in the dictionary increased over control in all cases, which is to be expected as augmentations to improve simplification and comprehension use more words (Fig. 3 (d)), and the Flesch Kincaid grade level largely decreased compared to control, especially for the simplified and to-do list augmentations, but augmentations like definitions increased in grade level (Fig. 3 (e)). This could be due to additional length and medical terminology in each of the definitions, which was also observed by some of the qualitative interview participants. Future work could explore further simplification augmentation of the definitions.
Quantitative Analysis
Figure 4: Results from our web-based survey for participants’ action understanding score for the Control, All, and Select conditions. We show the scores for all notes (a), synthetic and real notes (b), and scores across multiple subgroups (c). Generally, Select (definitions and simplification augmentations) leads to the best action understanding score and improvement over control.
In the survey study(n=200), we found that participants reading notes with either Select or All augmentations had statistically significantly higher action understanding scores, self-reported comprehension scores, and self-reported confidence scores when considering both real and synthetic note conditions in aggregate. Notably, with All augmentations compared to Control (no augmentations), action understanding scores increased from 0.537 to 0.605 (p=0.02), self-comprehension scores increased from 1.45 to 2.25 (p$<$ 0.001), and self-confidence scores increased from 2.17 to 2.95 (p$<$ 0.001) (Fig. 4 (a)). We find similar observations when comparing Select to Control.
We find that the trend of Select and All augmentations increase both self-reported and action understanding scores holds when looking at specific subgroups of the study population (Fig. 4 (c)). We note that Select and All augmentations had higher action understanding, self-comprehension, and self-confidence scores in each of the four subgroups.
When we split the participants based on the condition assigned of observing either synthetic notes or real notes, we found that Select and All augmentations statistically significantly improved action understanding scores only for the real notes and not for the synthetic notes (Fig. 4 (b)). For participants in the real notes condition, action understanding scores increased from 0.451 in Control to 0.541 in All augmentations (p=0.04); for participants in the synthetic notes condition, scores changed from 0.622 in Control to 0.669 in All (p=0.30).
The top two modifications that participants indicated they would use if they were to read another clinical note were simplifications chosen by 83.5% (95% CI 77.7-88.0) of participants and definitions chosen by 72.5% (95% CI 65.9-78.2) of participants. The remaining preferred modifications in order are to-do-list 42.0% (95% CI 35.4-48.9), FAQ 41.0% (95% CI 34.4-47.9), and key information 35.5 (95% CI 29.2-42.3). In the answers to the question “What did you use to answer these questions? Was it easy or hard?” the most common word was ”simplified” (occurred 97 times), followed by ”notes” (71), ”easy” (58), ”original” (54), and ”definitions” (54). This highlights that simplified and definitions augmentations were most useful to participants.
Qualitative Analysis
Participants (n=7) described three major themes around note-reading (Table F): (1) gathering information, (2) checking the narrative, and (3) understanding medical terminology. Participants shared that they read notes for facts such as lab results and future appointments, and to help their own memory and ensure they understood the visit correctly. Participants reported that they read notes in order to assess the narrative, for example ensuring that the clinician understood them and was aligned in their priorities. Many also described notes as a useful window into the clinician’s thinking: “You kind of want to have that insight into what the doctor’s take was on it” [P2]. All participants reported that medical terminology makes notes hard to understand, which can lead to time-consuming research and fear around the unknown.
All participants rated the definitions augmentation highest in the Likert ratings in terms of greatest benefit to note-reading. Participants described two main benefits with augmentations: (1) improved understanding of the note and (2) ease of use (Table F). Participants reported that augmentations made notes easier to read, less stressful, more personable, and in some cases prompted new questions and ideas. A few participants commented that the tool performed intuitively, highlighting jargon they would have struggled with and emphasizing takeaways they also found to be important. Many reported that it was convenient to have the tool integrated with the note and that the tool kept them focused: “When I didn’t have this, I would go off on websites that probably didn’t even apply to my type of cancer and they were doom and gloom, I would literally sit at my desktop and cry. It was so scary” [P4]. Participants also liked that they were able to toggle between the original note and the augmented note in order to check for accuracy.
Two major themes emerged around concerns: (1) That the augmentations may miss something or mislead patients and (2) more information could be confusing (Table F). One participant said, “I feel like there’s a tiny bit of anxiety being like, ’I’m having to trust that the AI got it right and that this is correct”’ [P2]. Others reported that augmentations may lose nuance in their oversimplification, misrepresent the gravity of the situation, or cause patients to think they understand more than they do. Participants also pointed out perceived discrepancies, such as a medication in the generated to do list that they were no longer taking.
Discussion
In this mixed methods study, including manual and automated error analysis, quantitative surveys, and qualitative interviews, we demonstrate that an LLM can be used to improve understanding of clinical notes. Participants performed better on action-oriented content questions and reported self-perceived improvements in understanding and confidence when reading notes with LLM-generated augmentations.
Subgroup analysis of vulnerable populations—less educated and nonwhite participants—showed that the benefits persist. These groups have previously reported the most benefit from reading clinical notes [27], and leveraging LLMs to improve note readability may be an important tool for organizations’ diversity, equity, and inclusion efforts. In our qualitative evaluation, patients reported that different people prefer different amounts of information about their health, and those information needs can vary over time. Further prompt engineering can explore custom augmentations that are tailored to patient readiness, their education level, or even their preferred language [28].
Through clinical review of LLM augmentations, we found that more work is needed to improve the accuracy of LLMs before they are deployed as standalone patient-facing tools. First, errors were more common in augmentations for real donated notes compared to augmentations for the carefully-written synthetic notes, implying that real-world deployment would still be impacted by the quality of the original note. Furthermore, while misleading errors were more rare and the majority of errors were flagged because of phrasing that could be confusing or misleading, patients’ trust was still impacted in qualitative interviews, even for those who were enthusiastic about the use of AI in medicine. We note that in our web-based survey, we did not find an impact in action understanding score due to errors, likely because the errors only impacted the ability to answer the action-oriented questions in one note, and participants could still view the original note. Especially in the presence of errors, carefully presenting LLM-generated advice is extremely important to study for future deployment [29]. In this study, we begin to consider this by presenting both the original and augmented notes for patients to reference. Further study of these impacts, as well as fine-tuning to reduce errors, can be effective in creating tools that truly empower patients [30].
Note-based augmentations can only be as accurate as the underlying note, and LLMs may amplify existing documentation errors and biases [31, 32]. In our study, patients preferred augmentations that provided medical definitions and simplified existing documentation. In the interviews, several participants also described the value of viewing the original note and emphasized the importance of reading notes not only for factual information but also to understand the clinician’s perspective and check for alignment. Further research is needed to understand how the clinicians’ and patients’ voices can be preserved while alleviating documentation and improving communication and education.
There are limitations to this study. We used a mix of real and synthetic notes in order to expand our pool of available notes for augmentation, but the synthetic notes were likely less complex and variable than the real donated notes. Augmentations led to improved scores on the real notes only, and most of the identified errors were found in the real notes. Future experiments may seek to evaluate the patient experience prospectively on real notes only. Evaluating a larger, more diverse cohort of patients at a medical center may also help to minimize the selection bias present in our study; participants who volunteered for our study may have been more likely to view AI favorably.
LLMs hold tremendous promise in their potential to transform healthcare. Careful design and implementation will provide the opportunity to change the patient-clinician relationship for the better as we move into this new era of health technology.
Author Contributions
All authors contributed to writing and conceptualizing the study. In addition, authors made the following contributions. NM generated augmentations, developed the user interface and survey, and conducted qualitative interviews. EBK guided study design and conducted analyses of augmentations, survey results, and qualitative interview results. BL contributed to study design, wrote synthetic clinical notes, conducted clinical error analysis, and coded qualitative interview results. CO contributed to study design and qualitative interviews. M Asiedu contributed to error analysis and project conceptualization, including project scoping, augmentation selection, and tool design. HM contributed to survey study design and conducted survey results analysis. M Agrawal contributed to project conceptualization and error analysis. AB contributed to project conceptualization and augmentation generation. TU and CER contributed to study design and preparation for institutional ethical review, as well as clinical error analysis. IBR contributed to project conceptualization, including augmentation selection. MG contributed to study design and interpreting results. DS guided the full project, including study design and analyses.
Acknowledgments
We would like to thank Hunter Lang, Shannon Shen, Daniel Ajayi, and Katie Liu for their helpful discussions throughout the project. We are grateful to the MIT Center for Clinical and Translational Research for their generous help with our human participant experiment. Furthermore, EBK was funded by the MIT CSAIL METEOR Fellowship. NM was partially funded by the Mark Foundation for Cancer Research. M Asiedu was supported by the Schmidt Science Postdoctoral Fellowship. M Agrawal was supported by the MIT Deshpande Center and the MLA@CSAIL initiative. This project was also supported by the National Center for Advancing Translational Sciences, National Institutes of Health, Award Number UL1TR002544. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Supplementary Material
Appendix A Supplementary File 1: Further Details on Proposed Augmentations
In the following subsections, for each of our augmentations we use, we discuss the benefits of Open Notes that we seek to enhance with each augmentation, challenges of Open Notes that may be addressed with each augmentation, prior work on that augmentation in both machine learning and medical communities, and the prompts used to generate each augmentation.
Definitions
Benefits To Enhance: Increase sense of control, improve understanding, identify misunderstandings or missing information, especially for those in poor health, support care partners, relieve anxiety, increase trust [33, 34, 31, 35, 36, 37].
Challenges To Address: Difficulty in understanding, especially for those who are older, less educated, and in poorer health, clinician burnout, oncologists’ concern that augmentations might compromise note integrity or take longer to write [38, 36, 39, 40, 41, 42].
Prior Work: Definition augmentations have been considered in the past, but many involved manual steps or focused on other domains, such as medical research papers [43, 44, 45]. These types of augmentations have been shown to be helpful [46].
Prompts: Make a list of the medical words that appear in this clinical note. Write each word as it appears in the note without modification. Define each of these terms. Definitions should be 1-2 sentences in simple English.
Simplification
Benefits To Enhance: Increase sense of control, improve understanding, support care partners, relieve anxiety, increase trust [35, 36, 37].
Challenges To Address: Difficulty in understanding, especially for those who are older, less educated, and in poorer health, clinician burnout, oncologists’ concern that augmentations might compromise note integrity or take longer to write [38, 36, 39, 40, 41, 42].
Prior Work: Simplifying medical content has been a widely studied topic in computer science [12], and has shown promising results in user studies [47]. We build upon this work with further experimentation with large language models.
Prompts: Rewrite this clinical note in simple English. Do not remove names of medications or diagnoses.
FAQ
Benefits To Enhance: Better prepare for visits, better understand, support care partners, relieve anxiety, increase trust and reassurance [35, 36, 37].
Challenges To Address: Difficulty in understanding, especially for those who are older, less educated, and in poorer health, clinician burnout, oncologists’ concern that augmentations might compromise note integrity or take longer to write [38, 36, 39, 40, 41, 42].
Prior Work: Large Language Models have been developed and tested on medical question answering [48], and the model responses are often preferred [49]. We focus on the patient perspective.
Prompts: Make a list of 7 common questions a patient might ask their doctor after reading this clinical note. Answer the questions in 2-3 sentences using simple English.
Key Information & To-Do List
Benefits To Enhance: Improve medication and lifestyle changes adherence, confirm and remember next steps, recollect care plan, relieve anxiety, increase trust and reassurance [33, 36, 37].
Challenges To Address: Increased feelings of anxiety, difficulty extracting useful information, especially action items and medications [36, 34].
Prior Work: Medications and physician action items have been extracted from clinical notes [50, 48], but we focus on patients and their perspectives.
Prompts: Make a list of the medications mentioned in this clinical note. Make a list of the upcoming appointments explicitly mentioned in this clinical note. If the note does not mention appointments, say ‘No Appointments’. Make a todo list of tasks the patient is responsible for before their next appointment based on the clinical note. Tasks need to be specific and actionable. Tasks should not be recurring events. If there are no such tasks mentioned in the note, say ”No Tasks”.
Appendix B Supplementary File 2: IRB Determinations
Notification of Expedited Approval
09-26-2022
To: Sontag, David Alexander
From: Leigh Firn, Chair
Study Title: OpenNotes: Making Clinical Notes More Friendly
Protocol #: 2208000725
Protocol Status: Active - Open to Enrollment
Expiration Date: 09-25-2025
Data Classification Level High Risk
Expedited Category: 5
Risk Level: No greater than minimal risk
The above-referenced protocol was approved following Expedited Review by the Committee on the Use of Humans as Experimental Subjects (COUHES).
Notification of Amendment Approval
12-05-2022
To: Sontag, David Alexander
From: Leigh Firn, Chair
Study Title: OpenNotes: Making Clinical Notes More Friendly
Protocol #: 2208000725A001
Protocol Status: Active - Open to Enrollment
Expiration Date: 09-24-2025
Data Classification Level High Risk
Expedited Category: 45CFR46.110(b)(1)(ii)
Risk Level: No greater than minimal risk
The amendment to the above-referenced protocol was approved following Expedited Review by the Committee on the Use of Humans as Experimental Subjects (COUHES). This approval covers the following change(s)/modification(s):
– The recruitment has been updated to include medical clinics.
– The protocol has been updated to include additional social media postings.
– The protocol has been updated to include an Amazon gift card for the first 10 participants.
Notification of Amendment Approval
04-13-2023
To: Sontag, David Alexander
From: Leigh Firn, Chair
Study Title: OpenNotes: Making Clinical Notes More Friendly
Protocol #: 2208000725A002
Protocol Status: Active - Open to Enrollment
Expiration Date: 09-24-2025
Data Classification Level High Risk
Expedited Category: 45CFR46.110(b)(1)(ii)
Risk Level: No greater than minimal risk
The amendment to the above-referenced protocol was approved following Expedited Review by the Committee on the Use of Humans as Experimental Subjects (COUHES). This approval covers the following change(s)/modification(s):
– Summary of Changes: Additional study procedures.
Updated consent, onboarding and recruitment docs.
Previously recruited and future participants who have completed the first part of the study (by donating their clinical notes) will now be invited for a virtual interview. During this interview, participants will be shown AI-modified version of one or more of their clinical notes and asked to comment on them. Modifications may include simplification, summarization, information extraction, or adding contextual information. Interview may be recorded and lasts approximately 45 minutes. Participants will be compensated with a \$30 Amazon gift card.
Notification of Amendment Approval
05-08-2023
To: Sontag, David Alexander
From: Leigh Firn, Chair
Study Title: OpenNotes: Making Clinical Notes More Friendly
Protocol #: 2208000725A003
Protocol Status: Active - Open to Enrollment
Expiration Date: 09-24-2025
Data Classification Level High Risk
Expedited Category: 45CFR46.110(b)(1)(ii)
Risk Level: No greater than minimal risk
The amendment to the above-referenced protocol was approved following Expedited Review by the Committee on the Use of Humans as Experimental Subjects (COUHES). This approval covers the following change(s)/modification(s):
– Adding additional crowdsourcing platforms to the protocol, including Prolific.
– Requesting a waiver of documentation of informed consent for Prolific due to their privacy policies.
Appendix C Supplemental Figure 3: Full user interface for augmented clinical notes
Figure 5: Full user interface for augmented clinical notes. (A) The user is able to switch between their original note and a simplified version using tabs. (B) The user’s note (either the original or simplified depending on the selected option) is displayed with clear formatting. (C) Medical terms are highlighted in bold green and underlined to indicate that the user is able to click on them. When a user clicks on a medical term, the page will automatically scroll to the corresponding definition in (G) (D) A title briefly describes the content of the clinical note. (E) Users can explore a series of questions designed to enhance their understanding of the note. By clicking on a question, its corresponding answer will be revealed. (F) Tabs enable seamless navigation between different augmentations, providing a convenient way to switch between them. The figure does not depict the Key Info and To-Do List augmentations, which d85isplay medication and appointment data as well as action items, respectively. (G) A scrollable sidebar is utilized to present definitions of medical terms. (H) An options menu allows the user to toggle the augmentations they wish to enable for their note.
Appendix D Supplemental Figure 4: Examples of Augmentations
Figure 6: Examples of augmentations using a sample excerpt from a synthetic clinical note. This includes the augmentations of simplification, definitions, frequently asked questions, key information, and to-do list.
Appendix E Supplemental File 3: Interview Questions

### v2

@@ 5.1 Study Population
For our interview study, all seven participants had accessed a portal for patients previously, demonstrated adequate health literacy, and all but one participant agreed with the statement that AI can have a beneficial impact in medicine.
Table 1: Mixed methods study population breakdown by demographics, experience with patient portals and notes, health literacy, and artificial intelligence sentiment.
Survey (N=200) Interview (N=7)
Highest Schooling 8th grade or less 3 0
High school graduate or GED 55 2
2-year college degree graduate 35 0
4-year college degree graduate 80 3
Masters or doctoral degree graduate 27 2
Race White 158 6
Black/African American 24 0
Asian 5 1
Native American/Alaska Native 1 0
Other 4 0
More than 1 7 0
Hispanic, Latino, or Spanish origin Yes 13 0
No 187 7
Gender Female 196 7
Male 3 0
Other 1 0
Age 40-45 42 0
46-55 68 3
56-65 61 3
66-75 28 1
$>$76 1 0
Employment Status Employed for wages 81 2
Self-employed 50 1
Homemaker 10 0
Unemployed 14 1
Retired 30 1
Unable to work 15 2
Accessed Portal for Notes? Yes 122 7
No 78 0
# Notes Last Year None 24 0
One 25 1
2 or 3 41 2
4 or more 26 4
Don’t know 6 0
General Health Literacy [10] Adequate 167 7
Inadequate 33 0
“Artificial Intelligence can have a beneficial impact in medicine” Strongly Agree 57 2
Agree 100 4
Neither agree nor disagree 32 1
Disagree 4 0
Strongly Disagree 7 0
5.2 Error Analysis
Figure 4: Error analysis of GPT-4 augmentations. (a) (b) and (c) show the number of errors in different categories, as determined by clinical reviewers. In (a), we have types of errors, such as hallucinations. The most frequent error was when GPT-4 assumed a certain context, e.g., assuming that a definition is specific to breast cancer when it truly is not. In (b), this is grouped by augmented clinical notes, where P represents a participant’s donated note, and S represents a synthetic note. There are very few errors for synthetic notes and a great deal of variation in the number of errors in donated clinical notes. In (c), the number of errors is shown in each of the augmentations, with definitions containing the most errors. (d) and (e) are the result of automated analysis of GPT-4 augmentations. The error bars on subfigures (d) and (e) represent confidence intervals computed using t-scores, note that we omit from the analysis empty note augmentations. (d) contains the percentage of words that are found in the Carnegie Mellon Pronouncing Dictionary Corpus Reader. Control has the least, as there are potentially more acronyms and jargon. (e) shows the Flesch Kincaid grade level score, which is generally comparable or less than control in the augmentations, except for definitions. From interviews, we know that definitions still do contain jargon that can be confusing.
We report the results of the error analysis in Figure 4. We found that GPT-4 can make misleading errors in the context of breast cancer clinical note augmentations, and identified 82 total errors. In total, 28 (34%) errors were considered to be potentially harmful and misleading (Figure 4 (a)). However, minor errors were most common, especially assumed context (Figure 4 (a)). For example, a definition augmentation specifying that lower outer quadrant referred specifically to the breast, though the term can refer to a specific part of other tissues as well. Many of these errors appeared in the definitions augmentation, and errors were especially common in one note’s augmentations (Fig. 4 (b-c)). In particular, most errors occurred in augmentations to real notes, while only one error appeared in augmentations to the synthetic notes (Figure 4 (b)).
In the automated analysis, we found that the percentage of words in the dictionary increased over control in all augmentations. Notably, when comparing control versus the simplified augmentation, the percentage increased for all notes, and the results of a paired-sample t-test are ($t=-4.61$, $p=0.0007$) (Figure 4 (d)). Moreover, the Flesch-Kincaid grade level decreased compared to control for the simplified and to-do list augmentations, when comparing control versus the simplified augmentation, the results of a paired-sample t-test are ($t=2.24$, $p=0.047$). However, augmentations like definitions increased in grade level (Figure 4 (e)), which could be due to additional length and medical terminology in each of the definitions. This was also observed by some of the qualitative interview participants. Future work could explore further simplification of the definitions and evaluation on a larger number of notes.
5.3 Quantitative Analysis
Figure 5: Results from our web-based survey for participants’ action understanding score for the Control, All, and Select conditions. We show the scores for all notes (a), synthetic and real notes (b), and scores across multiple subgroups (c). Generally, Select (definitions and simplification augmentations) leads to the best action understanding score and improvement over control. We note that both the All and Select conditions significantly increase action understanding score on the aggregate.
In the survey study (N=200), we found that participants reading notes with either Select or All augmentations had statistically significantly higher action understanding scores, self-reported comprehension scores, and self-reported confidence scores when considering both real and synthetic note conditions in aggregate. Notably, with All augmentations compared to Control (no augmentations), action understanding scores increased from 0.537 to 0.605 (p=0.02), self-comprehension scores increased from 1.45 to 2.25 (p$<$ 0.001), and self-confidence scores increased from 2.17 to 2.95 (p$<$ 0.001) (Figure 5 (a)). We found similar observations when comparing Select to Control.
We found that the trend of Select and All augmentations increasing self-reported and action understanding scores held when looking at specific subgroups of the study population, which we selected prior to analysis of the results (Figure 5 (c)). We note that Select and All augmentations had higher action understanding, self-comprehension, and self-confidence scores on average in each of the four subgroups, though we did not evaluate statistical significance. We also looked at other possible subgroup categorizations post-hoc and observed similar trends.
When we split the participants based on the condition assigned of observing either synthetic notes or real notes, we found that Select and All augmentations statistically significantly improved action understanding scores only for the real notes and not for the synthetic notes (Figure 5 (b)). For participants in the real notes condition, action understanding scores increased from 0.451 in Control to 0.541 in All augmentations (p=0.04); for participants in the synthetic notes condition, scores changed from 0.622 in Control to 0.669 in All (p=0.30).
The top two modifications that participants indicated they would use if they were to read another clinical note were simplifications chosen by 83.5% (95% CI 77.7-88.0) of participants and definitions chosen by 72.5% (95% CI 65.9-78.2) of participants. The remaining preferred modifications in order are To-Do-List 42.0% (95% CI 35.4-48.9), FAQ 41.0% (95% CI 34.4-47.9), and Key Information 35.5% (95% CI 29.2-42.3). In the answers to the question “What did you use to answer these questions? Was it easy or hard?”, the most common word was ”simplified” (occurred 97 times), followed by ”notes” (71), ”easy” (58), ”original” (54), and ”definitions” (54). This highlights that simplified and definitions augmentations were most useful to participants.
5.4 Qualitative Analysis
Table 2: Common themes from the interview study organized into three categories: Current practices and attitudes towards note reading, Wins with the modifications, and Concerns with the modifications.
Current practices and attitudes towards note reading
Reading notes to gather information “I wanted to know the dates when that was diagnosed, and maybe when was the last time I did a test” [P7] “Figure out when my appointments are. I’ll look at blood work, you know, if anything’s weird.” [P5]
Reading notes to check the narrative “[To find out] what my doc actually thought as opposed to what he said to me” [P4] “You kind of want that insight into like what the doctor’s take was on it” [P2]
Medical terminology makes notes difficult to understand “Usually, I have to Google some words. Especially in the beginning when you’re first diagnosed, there’s a lot of initials and words that a lot of people don’t understand” [P1] “At the time, I looked them up but they’re hard to retain even. It’s just I don’t use them in everyday language” [P2] “I have to do a lot of Googling though to see what this means and what that means and whether it’s significant or not.” [P3]
Wins with the modifications
Improved understanding of the note “It gives me more comfort and information going, oh, that’s what they mean” [P6] “It gives you all the information and it’s actually less stressful to read. I think because it’s not as overwhelming.” [P2] “It’s definitely easier to understand and at least one of the questions I had, I got answered here” [P5]
Ease of use “This is like a personalized Wikipedia page for my diagnosis” [P3] “It’s just nice to have it all there if I want it. I don’t have to look up everything but I can look up what I want to.” [P2]
Concerns with the modifications
Missing or misleading information “I feel like this makes assumptions that can’t really be made” [P5] “I guess I feel like blood work should be in there.” [P6]
More information can be confusing “It made me feel a little bit more worried because they were things that I hadn’t conceived of yet” [P2] “That’s a bit overly heavy on medical jargon.” [P3] “If I’m going to have a medication list, I think I would want the medication list of what I’m on or what I will be having, not everything that came up.” [P5]
Interview participants (N=7) described three major themes around note-reading (Table 2): (1) gathering information, (2) checking the narrative, and (3) understanding medical terminology. Participants shared that they read notes for facts such as lab results and future appointments, and to help their own memory and ensure they understood the visit correctly. Participants reported that they read notes in order to assess the narrative, for example ensuring that the clinician understood them and was aligned in their priorities. Many also described notes as a useful window into the clinician’s thinking: “You kind of want to have that insight into what the doctor’s take was on it” [P2]. All participants reported that medical terminology makes notes hard to understand, which can lead to time-consuming research and fear of the unknown.
All participants rated the definitions augmentation highest in the Likert ratings in terms of greatest benefit to note-reading. Participants described two main benefits with augmentations: (1) improved understanding of the note and (2) ease of use (Table 2). Participants reported that augmentations made notes easier to read, less stressful, more personable, and in some cases prompted new questions and ideas. A few participants commented that the tool performed intuitively, highlighting jargon they would have struggled with and emphasizing takeaways they also found to be important. Many reported that it was convenient to have the tool integrated with the note and that the tool kept them focused: “When I didn’t have this, I would go off on websites that probably didn’t even apply to my type of cancer and they were doom and gloom, I would literally sit at my desktop and cry. It was so scary” [P4]. Participants also liked that they were able to toggle between the original note and the augmented note in order to check for accuracy.
Two major themes emerged around concerns: (1) That the augmentations may miss something or mislead patients and (2) more information could be confusing (Table 2). One participant said, “I feel like there’s a tiny bit of anxiety being like, ’I’m having to trust that the AI got it right and that this is correct”’ [P2]. Others reported that augmentations may lose nuance in their oversimplification, misrepresent the gravity of the situation, or cause patients to think they understand more than they do. Participants also pointed out perceived discrepancies, such as a medication in the generated to-do list that they were no longer taking. We present our full codebook for the interviews in Appendix C.
6 Discussion
In this mixed methods study, including manual and automated error analysis, quantitative surveys, and qualitative interviews, we demonstrated that an LLM can be used to improve understanding of clinical notes. Participants performed better on action-oriented content questions and reported self-perceived improvements in understanding and confidence when reading notes with LLM-generated augmentations.
Subgroup analysis of vulnerable populations—less educated and nonwhite participants—showed that the benefits persist. These groups have previously reported the most benefit from reading clinical notes [15], and leveraging LLMs to improve note readability may be an important tool for organizations’ diversity, equity, and inclusion efforts. In our qualitative evaluation, patients reported that different people prefer different amounts of information about their health, and those information needs can vary over time.
Through clinical review of LLM augmentations, we found and characterized error modes, making it clearer what work is needed to improve the accuracy of LLMs before they are deployed as standalone patient-facing tools. First, errors were more common in augmentations for real donated notes compared to augmentations for the carefully-written synthetic notes, suggesting that a real-world deployment would still be impacted by the quality of the original note. Furthermore, while hallucinations were more rare, patients’ trust in qualitative interviews was still impacted by the errors we identified, even for those who were enthusiastic about the use of AI in medicine. We note that in our web-based survey, we did not find an impact in action understanding score due to errors, likely because the identified errors only impacted the ability to answer the action-oriented questions in one note, and participants could still view the original note. Especially in the presence of errors, carefully presenting LLM-generated advice is extremely important to study for future deployment [16]. In this study, we began to consider this by presenting both the original and augmented notes for patients to reference.
Relatedly, note-based augmentations can only be as accurate as the underlying note, and LLMs may amplify existing documentation errors and biases [4, 1]. In our study, patients preferred augmentations that provided medical definitions and simplified existing documentation. In the interviews, several participants also described the value of viewing the original note and emphasized the importance of reading notes not only for factual information but also to understand the clinician’s perspective and check for alignment. In the survey, we similarly hypothesize that Select augmentations slightly outperformed All augmentations because the additional augmentations in All may have opened up more potential misunderstandings.
There are limitations to this study. We used a mix of real and synthetic notes in order to expand our pool of available notes for augmentation, but the synthetic notes were likely less complex and variable than the real donated notes. Augmentations led to improved scores on the real notes only, and most of the identified errors were found in the real notes. Future experiments may seek to evaluate the patient experience prospectively on real notes only. Evaluating a larger, more diverse cohort of patients at a medical center may also help to minimize the selection bias present in our study; participants who volunteered for our study may have been more likely to view AI favorably. In addition, further research is needed to improve LLMs for clinical use, particularly for tools that are used by patients; additional fine-tuning to reduce errors [30], may help create tools that truly empower patients. However, given that errors may never be fully mitigated, another important direction to explore is improved interface design to make it easier for patients to notice potentially misleading facts in-line. Additional prompt engineering may help create custom augmentations that are tailored to patient readiness, their education level, or even their preferred language [51]. Further research is needed to understand how clinicians’ and patients’ voices can be preserved while alleviating documentation and improving communication and education.
7 Conclusion
In this work, we presented an end-to-end tool that improves information accessibility by giving patients the ability to view their clinical notes with five different LLM-enabled augmentations. We used the tool to study the effect of these augmentations both qualitatively and quantitatively, aiming to understand both patients’ attitudes towards AI and their understanding of actionable health information. We were also able to understand patients’ prioritization and preferences for choices of augmentation alongside clinician ratings of errors across these same augmentations. This characterization can inform future design choices in information accessibility, across domains.
Large language models hold tremendous promise in their potential to transform healthcare and empower patients with their own data. Careful design and implementation could provide the opportunity to change the patient-clinician relationship for the better as we move into this new era of health technology. However, thoughtful deployment and continued participatory design is required to modulate trust between users and AI, particularly in high-stakes scenarios like healthcare.
Appendix A User Interface Overview
Figure 6: Direct screenshot of user interface for augmented clinical notes, with annotations: (A) The user is able to switch between their original note and a simplified version using tabs. (B) The user’s note (either the original or simplified depending on the selected option) is displayed. (C) Medical terms are highlighted in bold green and underlined to indicate that the user is able to click on them. When a user clicks on a medical term, the page will automatically scroll to the corresponding definition in (G) (D) A title briefly describes the content of the clinical note. (E) Users can explore a series of questions about the note. By clicking on a question, its corresponding answer will be revealed. (F) Tabs enable navigation between different augmentations. The figure does not depict the Key Info and To-Do List augmentations, which display medication and appointment data as well as action items, respectively. (G) A scrollable sidebar is utilized to present definitions of medical terms. (H) An options menu allows the user to toggle the augmentations they wish to enable for their note.
Figure 7: User interface concept that aligns with Fig. 3.
Appendix B Interview Questions